# ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG

### v2

@@
Abstract
Conversational AI agents use Retrieval Augmented Generation (RAG) to provide verifiable document-grounded responses to user inquiries. However, many natural questions do not have good answers: about 25% contain false assumptions Yu et al. (2023), and over 50% are ambiguous Min et al. (2020). RAG agents need high-quality data to improve their responses to confusing questions. This paper presents a novel guided hallucination-based method to efficiently generate a diverse set of borderline out-of-scope confusing questions for a given document corpus. We conduct an empirical comparative evaluation of several large language models as RAG agents to measure the accuracy of confusion detection and appropriate response generation. We contribute a benchmark dataset to the public domain.
ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG
Zhiyuan Pengthanks: Equal contribution. Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian11footnotemark: 1 Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
1 Introduction
@@
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents Lewis et al. (2020); Guu et al. (2020); Gao et al. (2024). Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response. RAG performance is evaluated on dimensions like the response accuracy, faithfulness, completeness, answer and context relevance, and style Es et al. (2024); Saad-Falcon et al. (2024); Liu et al. (2024); Zheng et al. (2023). Errors may stem from a document-to-question mismatch, hallucinating parts of the response, wrong understanding of the document or the question.
Many user questions have no direct answer: some may rely on a false premise, some may be ambiguous in what they ask, and some are out of scope for the knowledge base. A well-designed RAG system should detect such confusing questions and respond appropriately, either by clarifying the confusion or seeking help.
Dataset RAG Setting Source Category Cut-off-knowledge Annotation Method
AMBIGQA Min et al. (2020) Wiki Multi-answers Human
CREPE Yu et al. (2023) Reddit False Premise Human
FalseQA Hu et al. (2023)  False Premise Human
KG-FPQ Zhu et al. (2024) Wiki False Premise LLM
FAITH Yuan et al. (2024) Wiki False Premise LLM
FLUB Li et al. (2024a) Baidu Tieba Cunning Texts Human
ScopeQA News Out-of-scope Human+LLM
Table 1: ScopeQA vs Existing Datasets
@@
Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type. As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge but neglects out-of-scope queries, which lack answers in retrieved documents. Borderline out-of-scope questions can confuse LLMs into making up a hallucinated answer. This type of confusion is the focus of this paper.
@@
While queries in MS MARCO Nguyen et al. (2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cut-off. ScopeQA fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to minimize human annotation effort. Our contributions are as follows:
@@
  • We propose and implement a method to generate borderline out-of-scope questions for given documents efficiently.
@@
  • We validate our synthetic data by training out-of-scope question detectors.
  • We evaluate LLMs as RAG agents for detecting confusion and generating responses.
  • Our code and dataset are released at https://github.com/zhiyuanpeng/ScopeQA.git.
2 Related Work
Benchmark Datasets
@@ 2 Related Work
Most of the recent work on LLM responses to confusing questions focuses on stand-alone questions answered using LLM’s parametric knowledge without a context document. The questions may be naturally collected from online sources Li et al. (2024b); Yu et al. (2023), written by human annotators Hu et al. (2023), or synthetically generated Zhu et al. (2024); Yuan et al. (2024). A comparison of these released benchmark datasets is shown in Table 1.
Question Generation
@@
Zhu et al. (2024); Yuan et al. (2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object from Wikipedia, replacing the object with a similar but incorrect one, and filling in a template. Content-grounded (non-confusing) synthetic data generation is studied in Yehudai et al. (2024); Wang et al. (2023b); Lee et al. (2023). Given the grounding passages, they generate data entries by few-shot prompting with task-specific examples, then filter to ensure data quality, faithfulness, and diversity.
Mitigation
@@
is commonly achieved with in-context learning  Zhu et al. (2024); Li et al. (2024b), chain-of-thought (CoT)  Wei et al. (2022) and fine-tuning. Experiments show that fine-tuning outperforms the other two methods Hu et al. (2023). CoT yields inconsistent improvements, while in-context learning improves performance with more examples Li et al. (2024b). Yuan et al. (2024) identifies  1% of attention heads responsible for confusion by false-premise questions, showing a 20% performance gain by constraining these heads.  Kim et al. (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
3 Methods
We developed the ScopeQA data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions. We assume that the retrieved document is the best available match for the question and focus on the response generation step of the RAG pipeline. In this step, given a standard prompt $p$, an out-of-scope question $q$, and a retrieved document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly and hallucinate.
@@ 3 Methods
1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix F.1)
@@
5:     Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix F.2)
6:     In the claims list $c_{1},\ldots,c_{n}$, replace the missing claims with the claims recovered in above step.
@@
8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only out-of-scope claims (Appendix F.11)
@@
9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix F.10)
10:Call LLMq to filter out any questions answerable in the context of $d$ (Appendix F.3)
3.1 Data Generation
@@
Here we describe our method of generating out-of-scope questions for a given set of documents. We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e. a few paragraphs long). We prefer documents that are novel to the LLMs under evaluation, e.g. news articles published later than all pretraining cutoff dates, so that the LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Thus, by using the Newscatcher Bugara et al. (2020), we collected 200 news articles published in 2024, each containing more than 150 words, for each topic (Appendix E). Sentences were selected sequentially from the top until the word count exceeded 300. Given a set of such documents, we generate multiple out-of-scope questions about each document and some in-scope ones for comparison.
We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the out-of-scope questions. We prefer but do not require LLMq to be larger and stronger than the LLMr under evaluation. For each document $d$, we call LLMq to generate a list of factual claims, then iteratively replace subsets of claims with hallucinated claims while maintaining semantic similarity to $d$ using guided hallucination, see Algorithm 1. We also generate in-scope questions for each document $d$ by calling LLMq directly.
3.2 Evaluation
@@
We first assessed the quality of ScopeQA by sampling a subset of questions and manually verifying whether they were indeed out-of-scope or not (Section 3.3). Next, we evaluated how well various LLMr models could defuse the out-of-scope questions in ScopeQA. Given the large number of out-of-scope questions, manual verification of LLMr responses is not feasible. Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question. To validate the effectiveness of this defusion detection method, we compared its performance with human annotations on a sampled dataset (Section 3.3). Formally, we define our tasks as follows:
Question generation:
@@
Defusion detection:
How accurately can AutoDefuseEval detect if an LLMr response to a out-of-scope question defuses the confusion? (Section 3.3)
Out-of-scope response:
How frequently does an LLMr, when responding to an out-of-scope question using different prompting methods, generate a response that successfully defuses the confusion? (Section 4.1)
Out-of-scope detection:
@@
Appendix A has the steps we run for evaluating the above tasks. We run steps 1, 2, and 3 for “out-of-scope detection” and 1, 4, 5, 6, and 7 for “out-of-scope response”. Following the self-consistency method Wang et al. (2023a), we perform multiple LLM calls in steps  2 and  5 of Algorithm A, taking the majority vote to determine the final label.

### v3

# ELOQ: Resources for Enhancing LLM Detection of Out-of-Scope Questions

@@
Abstract.
Large Language Models (LLMs) are widely used in Conversational AI systems to generate responses to user inquiries. However, many natural questions lack well-defined answers. While existing studies primarily focus on question types such as false premises, they often overlook out-of-scope questions, where the provided document is semantically highly similar to the query but does not contain the required answer. In this paper, we propose a guided hallucination-based method to efficiently generate a diverse set of out-of-scope questions from a given document corpus. We then evaluate multiple LLMs based on their effectiveness in confusion detection and appropriate response generation. Furthermore, we introduce an improved method for detecting such out-of-scope questions, enhancing the reliability of LLM-based question-answering systems.
Large Language Models, Question Answering, Out-of-Scope Question
*Both authors contributed equally to this work.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processing
1. Introduction
@@
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents (Lewis et al., 2020; Guu et al., 2020; Gao et al., 2024). Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response. RAG performance is evaluated on dimensions such as response accuracy, faithfulness, completeness, answer and context relevance, and style (Es et al., 2024; Saad-Falcon et al., 2024; Liu et al., 2024; Zheng et al., 2023). Errors may stem from a document-to-question mismatch, hallucinated parts of the response, or misunderstanding of the document or the question.
Many user questions have no direct answer: some rely on a false premise, others are ambiguous in their intent, and some are out-of-scope for the knowledge base. Prior studies indicate that about 25% of natural questions contain false assumptions (Yu et al., 2023) and over 50% are ambiguous (Min et al., 2020a). Additionally, out-of-scope errors have been identified as the leading cause of enterprise AI assistants generating responses that appear convincing but are, in fact, incorrect (Maharaj et al., 2024). A well-designed RAG system should implement an exception-handling mechanism that detect confusing questions and responds by clarifying the confusion, retrieving additional documents, or seeking assistance.
Dataset Source Category Cutoff knowledge Annotation Method
AMBIGQA (Min et al., 2020a) Wiki Multi-answers Human
CREPE (Yu et al., 2023) Reddit False Premise Human
FalseQA (Hu et al., 2023)  False Premise Human
KG-FPQ (Zhu et al., 2024) Wiki False Premise LLM
FAITH (Yuan et al., 2024) Wiki False Premise LLM
FLUB (Li et al., 2024a) Baidu Tieba Cunning Texts Human
ELOQ News Out-of-scope Human+LLM
Table 1. ELOQ vs Existing Datasets
@@
Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type. As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge, but often neglects out-of-scope queries for which no answers can be found in retrieved documents. Out-of-scope questions can mislead LLMs into generating hallucinated answers that appear correct, but are incorrect, which represents the most dangerous type of error (Maharaj et al., 2024). Our focus is to mitigate this type of confusion.
@@
This paper proposes a framework for automatically generating out-of-scope questions and evaluates its effectiveness in training classifiers to detect out-of-scope situations. While queries in MS MARCO (Nguyen et al., 2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cutoff. ELOQ fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to reduce human annotation effort. Our contributions are as follows:
@@
  • We propose and implement an efficient method for generating synthetic out-of-scope questions from a given set of documents.
@@
  • We demonstrate the utility of our synthetic data in training out-of-scope exception detectors.
  • We conduct a comparative evaluation of multiple LLMs as RAG agents to measure their confusion detection accuracy.
  • Our code  [^1] https://github.com/zhiyuanpeng/ELOQ.git and dataset  [^2] https://huggingface.co/datasets/zhiyuanpeng/ELOQ are made publicly available.
2. Related Work
2.1. Benchmark Datasets
@@
Recent work on LLM responses to confusing questions primarily focuses on stand-alone questions answered using the LLM’s pretraining knowledge, without a supporting context document. These questions may be naturally collected from online sources (Li et al., 2024b; Yu et al., 2023), written by human annotators (Hu et al., 2023), or synthetically generated (Zhu et al., 2024; Yuan et al., 2024). Released benchmark datasets with confusing questions include KG-FPQ (Zhu et al., 2024), FalseQA (Hu et al., 2023), FLUB (Li et al., 2024b), and CREPE (Yu et al., 2023). Earlier research on ambiguous open-domain questions (Min et al., 2020b) introduced AmbigNQ, a dataset of naturally ambiguous questions. Table 1 presents a comparison of these benchmark datasets.
2.2. Question Generation
@@
(Zhu et al., 2024; Yuan et al., 2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of (subject, relation, object) using Wikipedia data, replacing the object with a similar but incorrect one, and then filling in a template to generate a false premise question. Content-grounded (non-confusing) synthetic data generation has been studied in (Yehudai et al., 2024; Wang et al., 2023a; Lee et al., 2023). Given the grounding passages, these methods generate data entries using few-shot prompting with task-specific examples, followed by filtering to ensure data quality, faithfulness, and diversity.
2.3. Mitigation
@@
Mitigation is commonly achieved through in-context learning  (Zhu et al., 2024; Li et al., 2024b), chain-of-thought (CoT)  (Wei et al., 2022) and fine-tuning. Experiments show that fine-tuning outperforms both in-context learning and CoT reasoning (Hu et al., 2023). CoT reasoning leads to inconsistent improvements, whereas in-context learning improves performance as the number of examples increase (Li et al., 2024b). Yuan et al. (2024) identifies approximately 1% of attention heads as responsible for confusion caused by false-premise questions and demonstrates a 20% performance gain by constraining these heads.  Kim et al. (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
@@
1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix B.1)
@@
5:    Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix B.2)
6:    In the claims list $c_{1},\ldots,c_{n}$, replace the missing claims with the claims recovered in the above step.
@@
8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only out-of-scope claims (Appendix B.11)
@@
9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix B.10)
10:Call LLMq to filter out any questions answerable in the context of $d$ (Appendix B.3)
3. Methods
We developed the ELOQ data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions for a given document. We assume that the user has limited knowledge of the document, resulting in asking questions that appear relevant but are actually out-of-scope and cannot be answered based on the document’s content. In this step, given a standard prompt $p$, an out-of-scope question $q$, and a relevant document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question – a process we refer to as defusion (or de-confusion) – rather than attempting to answer the question directly and risking hallucination.
Algorithm 2 Evaluation steps
1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix B.6,  B.7,  B.8)
2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix B.3)
3:Aggregate $m$ predictions by majority vote
4:if $q$ is generated as out-of-scope then
5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix B.4)
6:     Aggregate $m$ predictions by majority vote
7:end if
Attr Gold Silver
# of documents 45 1955
# of in-scope questions 103 8,949
# of out-of-scope questions 113 10,105
average words per document 280 276
average words per in-scope question 16 16
average words per out-of-scope question 16 16
Table 2. Statistics of ELOQ. Gold is sampled from crawled 2000 news and annotated by annotators (Section 3.4).
3.1. Data Collection
@@
We focus on the scenario where the domain knowledge is maintained separately from the LLM in a document database. Hence, we prefer documents that are novel to the evaluated LLMs, such as news articles published after all pretraining cutoff dates, ensuring that LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Instead, they must generate responses grounded in their understanding of the news content. Therefore, using the Newscatcher (Bugara et al., 2020), we collected 200 news articles published between Jan 1, 2024, and September 30, 2024, for each topic333sport, business, science, food, politics, travel, entertainment, music, news, tech. We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e. a few paragraphs long). Thus, we select news articles with more than 150 words and, for each, sequentially extract sentences from the beginning until the word count exceeds 300. Finally, we collected 2,000 news and sampled a subset for human annotation. We represent the human-annotated data as Gold and the remaining data as Silver. Table 2 presents the dataset statistics.
3.2. Question Generation
We introduce our guided hallucination method, designed to generate out-of-scope questions from a given document. The process consists of three main steps: claim extraction, hallucination injection, and question generation.
3.2.1. Claim Extraction
The first step involves extracting claims from the document $d$ to serve as the basis for generating out-of-scope questions. Since our goal is to generate questions that appear related to $d$ but are actually unanswerable using its content, we start by obtaining a structured set of claims from the document. To achieve this, we use a LLM denoted as LLMq, to extract a set of $n$ claims $c_{1},c_{2},…,c_{n}$ from $d$. These claims represent key factual statements or assertions made in the document. We extract the claims in batches to minimize the number of LLM calls and improve computational efficiency. Once the claims are extracted, we partition them into disjoint subsets $S_{1},S_{2},…,S_{k}$. A simple partitioning scheme is used, such as:
$S_{j}=\{c_{i}\mid i\mod 3=j-1\},\quad\text{for }j=1,2,3.$
This ensures that claims are distributed evenly among the subsets, facilitating an efficient iterative process in the next step. Partitioning the claims allows us to selectively manipulate certain claims while preserving the overall document structure.
3.2.2. Hallucination Injection
The second step is to generate hallucinated claims that are similar to the factual claims generated in Section 3.2.1 but contain altered or missing information. The process follows an iterative masking and recovery approach over multiple rounds (e.g. 3). For each round:
  • For all claims in one subset $S_{j}$, we replace them with placeholders “(missing)”.
  • The modified claims list (with missing claims) is then provided to LLMq without access to $d$. The model is prompted to recover the missing claims based on the remaining claims in the list.
  • The recovered claims are inserted back into the claims list, replacing the placeholders.
  • This process is repeated for all subsets $S_{1},…,S_{k}$ across several rounds.
Finally, we discard claims that are explicitly supported by $d$. This ensures that only truly out-of-scope claims remain in our dataset.
3.2.3. Question Generation
The final step converts the hallucinated claims into out-of-scope questions while ensuring they cannot be answered using the document. Specifically, for each hallucinated claim, LLMq generates a concise and precise question focusing on a key aspect of the claim. The question is designed to appear related to the document while being unanswerable based on its content. Once the questions are generated, we further filter them by ensuring they are truly unanswerable. This is achieved by prompting LLMq to verify whether the generated questions can be answered using $d$. Any question that remains answerable in the context of $d$ is discarded.
In addition to generating out-of-scope questions, we also prompt LLMq to generate in-scope questions for comparison. The complete question generation process, along with its prompts, is outlined in Algorithm 1.
1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
  — 87.50 94.59
2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
  — 91.67 90.00
3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
  — 98.61 100.00
Table 3. Agreement scores and the performance of our method on the ELOQ-Gold.
3.3. Evaluation
@@
We first assessed the quality of ELOQ by sampling a subset of questions (ELOQ-Gold) and manually verifying whether they were truly out-of-scope (Section 3.4). Next, we evaluated how well various LLMr models could defuse the out-of-scope questions in ELOQ. Given the large number of out-of-scope questions, manual verification of LLMr responses is not feasible. Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question. To validate AutoDefuseEval, we compared its performance with human annotations (ELOQ-Gold) on a sampled dataset (Section 3.4). Formally, we define our tasks as follows:
Question generation::
@@
Defusion detection::
How accurately can AutoDefuseEval detect whether an LLMr response to a out-of-scope question defuses the confusion? (Section 3.4)
Out-of-scope response::
How often does an LLMr successfully defuse an out-of-scope question using different prompting methods? (Section 4.1)
Out-of-scope detection::
@@
Algorithm 2 has the steps we run for evaluating the above tasks. We run steps 1, 2, and 3 for “out-of-scope detection” and 1, 4, 5, 6, and 7 for “out-of-scope response”. Following the self-consistency method (Wang et al., 2023b), we perform multiple LLM calls in steps  2 and  5 of Algorithm 2, taking the majority vote to determine the final label.
3.4. Human Annotation
Figure 1. Confusion matrix of out-of-scope and defusion on ELOQ-Gold.
Following the methodology from (Cochran, 1977), we sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators. As shown in Table 3, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is out-of-scope and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency. Two additional annotators manually resolved all disagreements to establish ground truth labels. All annotators were computer science graduate students recruited via email and incentivized with complimentary meals. They were trained by first reviewing the annotation guidelines and then labeling a small dataset previously annotated by the authors. The annotation guidelines are released along with the data and code.
We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of the generated labels. “Group’s Agree Acc” uses the instances where both annotators within the same group agreed, treating this consensus as the ground truth. This method yielded higher accuracy than “Annotator Acc,” leading us to resolve disagreements with two additional annotators, ultimately establishing the final labeled data as ground truth.
As reflected in “Ground Truth Acc,” in Table  3, our method achieved 94.91% accuracy in generating out-of-scope and in-scope questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion. As shown in Figure 1, the confusion matrix on the left aligns closely with the ground truth, with minimal errors (5 false negatives and 6 false positives). The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
Model Name Quantization Knowledge Cutoff Date
Llama 3.2 3B Instruct Turbo FP8 2023/12
Llama 3.1 8B Instruct Turbo FP8 2023/12
Llama 3.1 70B Instruct Turbo FP8 2023/12
Llama 3.3 70B Instruct Turbo FP8 2023/12
Mistral (7B) Instruct v0.3 FP16 2024/5
gpt-3.5-turbo   — 2021/9
Table 4. Different LLMr being evaluated. In ELOQ-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cutoff date of 5/22/2024.
4. Experimental Results